Grade 10Analytical Geometry

Two equal cones are touching each other completely at the base circle. Given that the distance between the two vertices is 16cm and the diameter of the base circle is 12cm find the total surface area of the solid. Bhai bhai bhai koi ans kardo

Vaibhav gupta
8 Years agoGrade 10

1 Answer

Piyush Bhujbal
8 Years ago
(Curved surface area)*2=total surface areaBoth are similar conesThere for we can deduce slant height which is equal to 10cm.CSA=πrl=π*6*20=120πTotal surface area=(CSA)*2=120π*2=240π